40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 25 13:23:45 2021, Anchal, Summary, BHD, Before photos of BSC 
    Reply  Mon Oct 25 17:37:42 2021, Anchal, Summary, BHD, Part I of BHR upgrade - Removed optics from BSC rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpgrn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg
       Reply  Tue Oct 26 16:56:22 2021, Paco, Summary, BHD, Part I of BHR upgrade - Locked PMC and IMC 
          Reply  Wed Oct 27 16:27:16 2021, Anchal, Summary, BHD, Part II of BHR upgrade - Prep OMCNorthBefore.jpegOMCNorthAfter.jpeg
             Reply  Wed Oct 27 16:31:35 2021, Anchal, Summary, BHD, Part III of BHR upgrade - Removal of PR2 Small Suspension 
                Reply  Thu Oct 28 16:32:32 2021, Paco, Summary, BHD, Part IV of BHR upgrade - Removal of BSC eastern optics 
                   Reply  Fri Nov 5 12:21:16 2021, Anchal, Summary, BHD, Part VI of BHR upgrade - Removal of ITMYC optics XendFlowBenchLeftEnd.jpgXendFlowBenchMiddle.jpgXendFlowBenchRightEnd.jpg
                      Reply  Wed Dec 15 16:11:38 2021, Anchal, Summary, BHD, Part VIII of BHR upgrade - Placed LO1 signal-2021-12-15-161437.jpeg
                         Reply  Wed Dec 15 17:44:18 2021, Koji, Summary, BHD, Part VIII of BHR upgrade - Placed LO1 
                            Reply  Thu Jan 6 21:04:41 2022, Anchal, Summary, BHD, Part VIII of BHR upgrade - LO1 OSEMs inserted 
                               Reply  Mon Jan 10 14:52:51 2022, Anchal, Summary, BHD, LO1 OSEMs roughly calibrated and noise measured LO1_vs_ETMX_OSEM_Spectrum_LF_x30_Gain.pdfLO1_vs_ETMX_OSEM_Spectrum_HF_x30_Gain.pdf
                                  Reply  Mon Jan 10 18:36:41 2022, Anchal, Summary, BHD, LO1 free swinging test set to trigger 
                                     Reply  Tue Jan 11 12:19:12 2022, Anchal, Summary, BHD, LO1 Input Matrix Diagonalization performed. LO1_SUS_InpMat_Diagnolization.pdfLO1_FreeSwingData_PeakFitting.pdf
                               Reply  Mon Jan 24 18:53:09 2022, Anchal, Summary, BHD, Part VIII of BHR upgrade - LO2 placed and OSEMs tuned 
                                  Reply  Tue Jan 25 10:55:02 2022, Anchal, Summary, BHD, Part VIII of BHR upgrade - LO2 input matrix diagonalization performed. 
                                     Reply  Tue Jan 25 16:42:03 2022, Anchal, Summary, BHD, Reduced filter gains in all damped new SOS 
                         Reply  Thu Jan 6 11:54:20 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - Placed AS1 and AS4 
                            Reply  Fri Jan 7 16:17:42 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - Placed AS1 and AS4 filters 
                               Reply  Sat Jan 8 16:01:42 2022, Paco, Summary, BHD, Part IX of BHR upgrade - Placed LO2 filters 
                                  Reply  Tue Jan 18 14:26:09 2022, Paco, Summary, BHD, Part IX of BHR upgrade - Placed remaining filters SR2, PR3, PR2 
                               Reply  Tue Jan 11 18:18:29 2022, Anchal, Summary, BHD, Attempted OSEM installation on AS1 
                                  Reply  Tue Jan 11 18:40:25 2022, Anchal, Summary, BHD, AS1 Sat Amp has a PCB issue signal-2022-01-11-184917.jpeg
                                     Reply  Thu Jan 13 09:48:41 2022, Anchal, Summary, BHD, AS1 Sat Amp fixed signal-2022-01-13-094823_001.jpegsignal-2022-01-13-094823_002.jpeg
                                  Reply  Thu Jan 13 12:24:08 2022, Anchal, Summary, BHD, AS1 SD and LR magnets broke 
                                     Reply  Mon Jan 24 21:18:40 2022, Paco, Summary, BHD, Part IX of BHR upgrade - AS1 placed and OSEM tuned 
                                        Reply  Tue Jan 25 11:02:54 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - AS1 free swing test failed AS1_freeSwingTestFailed.pdf
                                           Reply  Thu Jan 27 08:27:34 2022, Paco, Summary, BHD, Part IX of BHR upgrade - AS1 inspection and correction 
                                              Reply  Wed Feb 2 16:32:02 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - AS1 free swing test still not good AS1_FreeSwingData_PeakFitting_20220125.pdfAS1_SUS_InpMat_Diagnolization_20220202.pdf
                                                 Reply  Tue Mar 15 18:42:37 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - AS1 resuspended and OSEMs tuned 
                                                    Reply  Tue Mar 15 18:45:12 2022, Anchal, Summary, BHD, Part X of BHR upgrade - BHDBS Path setup 
                            Reply  Thu Jan 13 12:29:27 2022, Anchal, Summary, BHD, AS4 LR magnet broke 
                               Reply  Fri Jan 14 18:12:47 2022, Anchal, Summary, BHD, AS4 placed in ITMY Chamber, OSEMs connected 
                                  Reply  Tue Jan 18 18:19:22 2022, Koji, Summary, BHD, AS4 placed in ITMY Chamber, OSEMs connected Screen_Shot_2022-01-18_at_17.48.16.pngcable_swap.png
                                     Reply  Wed Jan 19 12:56:52 2022, Paco, Summary, BHD, AS4 OSEMs installation 
                                        Reply  Thu Jan 20 16:42:55 2022, Paco, Update, BHD, AS4 OSEMs installation - part 2 
                                           Reply  Thu Jan 20 18:16:29 2022, Anchal, Update, BHD, AS4 set to trigger free swing test 
                                              Reply  Fri Jan 21 16:40:10 2022, Anchal, Update, BHD, AS4 Input Matrix Diagonalization performed.  AS4_SUS_InpMat_Diagnolization_20220121.pdfAS4_FreeSwingData_PeakFitting_20220121.pdf
                                                 Reply  Mon Jan 24 17:12:27 2022, rana, Update, BHD, AS4 Input Matrix Diagonalization performed.  
                                                    Reply  Mon Jan 24 20:48:48 2022, Anchal, Update, BHD, AS4 Input Matrix Diagonalization performed.  
                   Reply  Tue Jan 18 15:19:49 2022, Paco, Summary, BHD, Part IV of BHR upgrade - Replaced old SR3 with SOS SR2. OSEM tuning attemtped. PXL_20220119_012954365.jpgPXL_20220119_013003522.jpg
                      Reply  Tue Jan 18 18:16:28 2022, Anchal, Summary, BHD, Part IV of BHR upgrade - SR2 Sat Amp Box inspection 
                         Reply  Wed Jan 19 12:50:10 2022, Anchal, Summary, BHD, Part IV of BHR upgrade - SR2 OSEM tuning progress. 
                            Reply  Thu Jan 20 17:03:36 2022, Anchal, Summary, BHD, Part IV of BHR upgrade - SR2 OSEM tuned. DJI_0144.JPG
                               Reply  Thu Jan 20 18:41:55 2022, Anchal, Summary, BHD, SR2 set to trigger free swing test 
                                  Reply  Fri Jan 21 11:24:42 2022, Anchal, Summary, BHD, SR2 Input Matrix Diagonalization performed.  SR2_SUS_InpMat_Diagnolization_20220121.pdfSR2_FreeSwingData_PeakFitting_20220121.pdf
                Reply  Thu Oct 28 17:01:54 2021, Anchal, Summary, BHD, Part III of BHR upgrade - Adding temp fixed flat mirror for PR2 
                Reply  Thu Jan 27 17:57:35 2022, Anchal, Summary, BHD, Part III of BHR upgrade - Replaced small suspended PR3 with new SOS PR3 and OSEM tuning 
                   Reply  Thu Jan 27 18:03:36 2022, Paco, Summary, BHD, Part III of BHR upgrade - Install PR2, balance, and attempt OSEM tuning 
                      Reply  Thu Jan 27 20:46:38 2022, Koji, Summary, BHD, Part III of BHR upgrade - Install PR2, balance, and attempt OSEM tuning Screen_Shot_2022-01-27_at_20.33.56.pngScreen_Shot_2022-01-27_at_21.30.16.png
                         Reply  Fri Jan 28 10:37:59 2022, Paco, Summary, BHD, Part III of BHR upgrade - PR2 OSEM finalized, reinsall LO1 OSEMs 
                            Reply  Fri Jan 28 11:30:52 2022, Koji, Summary, BHD, Part III of BHR upgrade - PR2 OSEM finalized, reinsall LO1 OSEMs 
                               Reply  Fri Jan 28 16:35:18 2022, Anchal, Summary, BHD, AS1, PR2 and PR3 set to trigger free swing test 
                            Reply  Wed Feb 2 16:22:00 2022, Anchal, Summary, BHD, Part III of BHR upgrade - PR2 inpute matrix diagonalized PR2_FreeSwingData_PeakFitting_20220125.pdfPR2_SUS_InpMat_Diagnolization_20220202.pdf
                   Reply  Wed Feb 2 16:27:57 2022, Anchal, Summary, BHD, Part III of BHR upgrade - PR3 inpute matrix diagonalized PR3_FreeSwingData_PeakFitting_20220125.pdfPR3_SUS_InpMat_Diagnolization_20220202.pdf
             Reply  Wed Oct 27 16:38:02 2021, Anchal, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
                Reply  Wed Oct 27 18:11:37 2021, Koji, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
                Reply  Wed Oct 27 18:16:45 2021, Koji, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
             Reply  Thu Mar 17 17:23:22 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - Green laser alignment 
                Reply  Mon Mar 21 18:24:07 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - Green laser alignment on Yarm 
                   Reply  Thu Mar 24 19:20:28 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - Green laser alignment on Yarm 
                      Reply  Fri Mar 25 18:20:28 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - Green laser alignment on Yarm 
                         Reply  Sat Mar 26 13:04:38 2022, Paco, Summary, BHD, Part IIa of BHR upgrade - Green laser alignment on Yarm 
                            Reply  Thu Mar 31 18:58:16 2022, Ian MacMillan, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Xarm 
                               Reply  Tue Apr 5 18:15:06 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Xarm 
                                  Reply  Wed Apr 6 01:20:48 2022, Koji, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Xarm PXL_20220406_060846581.jpgPXL_20220406_014205726.jpgPXL_20220406_060324812.jpgPXL_20220406_015259326.jpg
                                     Reply  Wed Apr 6 22:51:47 2022, Koji, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Xarm PXL_20220406_213356237.MP.jpgPXL_20220406_212936502.MP.jpg
                                  Reply  Wed Apr 6 12:03:51 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Yarm 
                                     Reply  Thu Apr 7 17:59:51 2022, Anchal, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Yarm 
                                        Reply  Fri Apr 8 17:21:31 2022, Ian MacMillan, Summary, BHD, Part IIa of BHR upgrade - IR laser alignment on Yarm 
                                           Reply  Mon Apr 11 21:13:21 2022, Paco, Summary, BHD, Part IIa of BHR upgrade - POY11 debugging 
                                              Reply  Mon Apr 11 21:44:14 2022, Koji, Summary, BHD, Part IIa of BHR upgrade - POY11 debugging D990543A1.PDFPXL_20220412_023941730.jpgPXL_20220412_023950567.jpg
                                                 Reply  Thu Apr 14 11:52:57 2022, yehonathan, Summary, BHD, Part IIa of BHR upgrade - POY11 debugging Screenshot_2022-04-14_15-53-39.pngPXL_20220414_220616875.jpg
                                                    Reply  Thu Apr 14 18:39:13 2022, Paco, Summary, BHD, Part IIa of BHR upgrade - POY11 debugging 
                                                       Reply  Mon Apr 18 17:53:45 2022, Paco, Summary, BHD, Part IIa of BHR upgrade - POY11 debugging 
Message ID: 16770     Entry time: Mon Apr 11 21:13:21 2022     In reply to: 16768     Reply to this: 16771
Author: Paco 
Type: Summary 
Category: BHD 
Subject: Part IIa of BHR upgrade - POY11 debugging 

[Paco, Koji]

I asked Koji for some advice regarding closing the loop on YARM using POY11. A few things seemed off including

  • The YARM transmission; which was peaking at ~ 20 (typically, TRY is normalized so that under nominal input power conditions we see TRY in the range [0, 1])
  • The POY11 DCPD level was quite low; we expect a few tens of uW in this low power configuration where no more than 100 mW are going through IMC.

We looked at the POY11 RFPD first. We tried flashing an incandescent lamp in-situ and saw some weak response using an oscilloscope and the DC Out readout. We then used the OPHIR power meter and recorded ~ 1.2 uW of light incident on the POY11 RFPD... Initially, we suspected our beam was not filling the PD sensitive area (~ 2 mm diameter), but a quick estimate using the 200 mm focusing lens currently installed in the ITMY table gave us quite a generous margin of error... so we questioned the OPHIR measurement. We swapped the power meter and this time got ~ 18.4 uW, which is more in line with what we expected (phew).

Moving on, from the POY11 RFPD responsivity, and our ~ 20 uW of incident power, we expected on the order of a 20 mV of DC Output, but weren't really seeing this on the scope, so we decided to test the pins on the power of the RFPD. The DB15 cable not only supplies bipolar 15 VDC but also monitors several other test points such as Tsens, or DCout in the RFPD. We quickly noticed a weird signal on the ENAB testpoint, so we removed POY11 RFPD from the ITMY table and took it to the PD testbench. After redoing the soldering on the breakout board and RF amplifier (ZXX-500LN+), which we tested separately, we saw the expected behavior using a +- 15 VDC power supply... thus verifying that the RFPD and breakout board seemed to work ok. We turned our attention to the upstream DB15 connection, and after quickly checking the newly run cables, we ended up debugging the eurocrate PD interface. After attempting a simple power cycle and failing, we removed this card and looked at the schematic. It would seem that the logic enabling ICs (one or both) failed, thus preventing the card from enabling its outputs correctly.... We bypassed the logic by soldering pins 4,8,14 on U1, U20 and then checked the circuit in-situ, and we saw it worked fine again.

Now none of the status LEDs (which are driven by the logic IC portion) work on this card, but the card itself works fine at least for POY11.

We moved on, and installed the DB15 cables, checking the functionality at every step... Then we looked at the POY11_I_ERR signal and were happy to see nice pdh wavelets. We pushed forward a little bit more to try and lock YARM. First, we went to the Y end and centered the ETMY Oplev so as to register the position where the YARM is flashing... The ITMY Oplev is still not online. Then, we optimized the ETMY damping gains somewhat to try and make it less noisy, and finally, played with the LSC YARM loop gain to attempt locking. This last push was not as successful, but we have an idea of what next steps are needed to reduce the SUS noise, including

  • Install ITMY Oplevs, and close loop
  • Optimize ITMY damping gains

To be continued....

ELOG V3.1.3-